### v1

# TumorPred: A Computational Framework Implemented via an R/Shiny Web Application for Parameter Estimation and Sensitivity Analysis in Compartmental Brain Modeling

Abstract
It is difficult or infeasible to directly measure how much of a drug actually enters the human brain and a brain tumor, how long it remains there, and to estimate drug-specific or patient-specific parameters, as well as how changes in these parameters influence model outputs and pharmacokinetic characteristics. Compartmental modeling offers a powerful mathematical approach to describe drug distribution and elimination in the body using systems of differential equations. This study introduces TumorPred, an R/Shiny-based web application designed for model simulation, sensitivity analysis, and pharmacokinetic parameter calculation in a permeability-limited four-compartment brain model. The model closely mimics human brain functionality for drug delivery and aims to predict the pharmacokinetics of drugs in the brain blood, brain mass, and cranial and spinal cerebrospinal fluid (CSF) of the human brain. The app provides real-time output updates in response to input modifications and allows users to visualize and download simulated plots and data tables. The computational accuracy of TumorPred is validated against results from the Simcyp Simulator (Certara Inc.). TumorPred is freely accessible and serves as an invaluable computational tool and data-driven resource for advancing drug development and optimizing treatment strategies for more effective brain cancer therapy.
Keywords
R/Shiny; pharmacokinetic modeling; brain cancer; parameter estimation; sensitivity analysis
Code metadata
Nr. Code metadata description
C1 Current code version 1.0
C2 Permanent link to code/repository used for this code version https://github.com/CharukaWick/TumorPred_Version_1.0
C3 Permanent link to Reproducible Capsule https://github.com/CharukaWick/TumorPred_Version_1.0
C4 Legal Code License MIT License
C5 Code versioning system used git
C6 Software code languages, tools, and services used R
C7 Compilation requirements, operating environments & dependencies None
C8 If available Link to developer documentation/manual N/A
C9 Support email for questions gi6036@wayne.edu
1.   Introduction
Quantitative knowledge of drug absorption, distribution, metabolism, and excretion (ADME) of many new and existing drugs in the human central nerve system (CNS) is essential for the development of new drugs and optimal use of existing drugs for brain cancer. However, direct measurement of drug penetration in the human brain and brain tumors is difficult or infeasible given the challenge of sampling and limitation of currently available imaging and analytical technologies. Physiologically based pharmacokinetic (PBPK) modeling offers a mechanism based computational approach for quantitatively predicting the ADME of drugs, as it can incorporate both biological system-specific and drug-specific data into a pharmacokinetic model. These PBPK models consist of complex systems of coupled ordinary differential equations. As a result, there is growing interest in developing user-friendly computational tools to solve such models. These tools enable multidisciplinary teams in the pharmaceutical industry, including researchers and clinicians without programming expertise, to effectively support decision-making in drug development.
To this end, we developed a computational framework (TumorPred) via an R/Shiny web application. TumorPred offers three main functionalities. Module 1 for model simulations helps direct calculation of drug concentration profiles and estimation of pharmacokinetic parameters. Module 2 for sensitivity analysis is designed to understand which parameters are the most influential in predicting drug concentration profiles in the CNS. Module 3 for parameter estimation helps in estimating drug-specific and patient-specific parameters which is crucial for modeling drug behavior in the body and ensuring the effectiveness and safety of medications, while inaccurate estimates can weaken pharmacokinetic predictions, slowing drug development and clinical practices.
We incorporated a 4-compartment permeability-limited brain model introduced in [1] into the TumorPred app, as this model is considered one of the gold standards for predicting drug penetration, allowing a wide range of users to benefit from the TumorPred. As supporting evidence, the same brain model is implemented in the Simcyp Simulator[2], a widely used commercial software for drug discovery. The novelty of TumorPred lies in its ability to estimate both drug-specific and system-specific model parameters using a parameter estimation algorithm, rather than relying on values from the literature or in vitro–in vivo studies. The app is programmed in R and utilizes Shiny as the web application framework, taking advantages of key R packages.
2.   Software description
We developed R/Shiny web applications using RStudio, one of the most widely adopted integrated development environments (IDEs) for R programming. The application code is organized in a single file containing three main components: the user interface (UI), the server, and the shinyApp function, as illustrated in Figure 1(d). In our design, the graphical user interface (GUI) is divided into a sidebar panel(left) and a main panel(right) as shown in Figure 1(a). The UI defines the graphical layout of the application. Inputs provided by the user through the UI are transferred to the server, where they are processed. The command shinyApp(ui = ui, server = server) links the UI and server into a single interactive web application. The server then generates outputs such as figures, tables, or other graphical elements which are displayed in the positions specified by the UI. Furthermore, a www directory (Figure 1(b))can be created in the root folder of a Shiny application to store external resources (e.g., scripts, images, excel files, or videos) that can be directly referenced within the app. The system of differential equations that defined 4-CNS brain model is shown in Figure 1(c). The description of the parameters and the reference values can be found in the supplementary materials. Finally, the TumorPred application has been deployed on shinyapps.io, making it accessible to a broader scientific and user community.
Figure 1: The structure of the TumorPred software.
 2.1  Input file preparation
Drug penetration into and distribution within the four CNS compartments are driven by the plasma drug concentration–time profile. We assume user has a set of state vector observations available for parameter estimation. Therefore, the user input file (.csv) must contain the following data columns: time, plasma, observed data for each compartment, and optionally, a set of parameter values (if the user wishes to perform a sensitivity analysis independently of parameter estimation). A sample input file can be accessed through the GUI of TumorPred; it contains data gathered from the Simcyp Simulator following the oral administration of a single 10 mg dose of abemaciclib drug. Note that the user must retain the original column names in the input file, as the server relies on them to read and process the data upon submission.
 2.2  Model simulations
Module 1 allows users to perform model simulations. The results include an estimated concentration data table, concentration-time plots, and key pharmacokinetic parameters (Cmax, Tmax, and AUC), all of which can be viewed and downloaded. For demonstration purposes, a sample input file is available through the GUI; this sample can be used in place of the user’s own data.
 2.3  Sensitivity analysis
Sensitivity analysis is a crucial method for identifying the parameters that most significantly influence the outcomes of a model. In the context of this study, sensitivity analysis can be approached in two ways. First, following a rigorous parameter estimation process, users can update the input file with the estimated parameters to proceed with the sensitivity analysis. Alternatively, if the user already has all the parameter values for a specific drug under investigation, the sensitivity analysis can be conducted directly without the need for parameter estimation. For example, a sensitivity analysis conducted using TumorPred (Figure 2) reveals that the drug concentration in the brain mass (Cbm) is highly sensitive to the parameter PSB, unlike Cbb, Cccsf, and Cscsf, which remain unaffected.
Figure 2: Sensitivity analysis results for the parameter PSB.
 2.4  Parameter estimation
To find the best-fit parameter values for a given drug and patient population, we define a loss function ($\text{loss}=\min_{p}\sum_{i=1}^{n}\sum_{j=1}^{k}\frac{(\hat{y}_{j}(t_{i})-y_{j}(t_{i},p))^{2}}{2\sigma^{2}_{j}}$) that minimizes the sum of the squares of the differences between the observed data points and the values predicted by the model such that $y(t,p)$ satisfies the system of ODEs. Here, $n$ is the number of observation points, $k$ is the number of equations in the ODE system, $\hat{y}_{j}(t_{i})$ denotes the observed data, $y_{j}(t_{i},p)$ denotes the model-predicted values, and $\sigma^{2}_{j}$ is the variance of the observed data for component $j$. We then pass this loss function to the Differential Evolution (DE) algorithm, a global optimization method available in the DEoptim library in R. Within the loss function, the system of differential equations is solved using the Livermore Solver for Ordinary Differential Equations (LSODA), which features automatic method switching. This allows the solver to automatically determine whether the system is stiff and choose the appropriate integration method, removing the need for the user to make that determination.
Figure 3: TumorPred outputs under parameter estimation.
TumorPred is primarily designed to estimate all 27 parameters of the 4-CNS brain model. However, estimating all 27 parameters simultaneously can be computationally expensive. The solution may sometimes fail to converge due to several factors: the high stiffness of the system of ODEs, the high sensitivity of drug-specific and patient-specific parameters, numerical errors from finite computer precision, or a lack of high-quality data. Furthermore, the computational time depends on the number of iterations, the relative tolerance, and the value-to-reach (VTR) termination threshold for the loss function. Values for some drug- and patient-specific parameters can be obtained from experimental data or literature. In such cases, these values can be entered as the minimum and maximum bounds for the parameters via the GUI, which reduces the algorithm’s computational burden. Following a successful simulation, the user can view and download a list of the estimated model parameters, the corresponding concentration-time profiles, log concentration-time profiles, and key pharmacokinetic parameters (Cmax, Tmax, and AUC), as shown in Figure 3.
3.   Software validation
For validation and simplicity, we selected six drug- and patient-specific parameters ($V_{bb}$, $V_{bm}$, $V_{ccsf}$, $V_{scsf}$, $fu_{bb}$, and $\lambda_{ccsf}$) to estimate. Table 1 presents the parameter estimation results, including the reference values (Ref.), the estimated values (Est.), and the absolute errors ($\mathbf{Error}=|\mathbf{Ref}-\mathbf{Est}|$). A figure depicting the convergence of each parameter throughout the iterations for this simulation is provided in the supplementary materials. Upon successful estimation of the parameters, we used the TumorPred app to estimate drug concentration profiles and compared these results with the data obtained from the Simcyp simulator, as shown in Figure 4. Both Table 1 and Figure 4 show that the estimated parameters are in strong agreement with the reference values, and the estimated drug concentration profiles closely match the data from the Simcyp simulator. This guarantees the accuracy of the TumorPred calculations and suggests that TumorPred is a promising tool for drug development and optimizing treatment strategies for more effective brain cancer therapy.
Table 1: Parameter comparison: reference vs. predicted values with absolute errors.
Parameter. $V_{bb}$ $V_{bm}$ $V_{ccsf}$ $V_{scsf}$ $fu_{bb}$ $\lambda_{ccsf}$
Ref. 0.064952435 1.104115461 0.103984624 0.025996156 0.125000000 0.026000000
Est. 0.064952243 1.014766461 0.103953204 0.025996112 0.124999861 0.025691588
Error $1.92\times 10^{-7}$ $8.93\times 10^{-2}$ $3.14\times 10^{-5}$ $4.40\times 10^{-8}$ $1.39\times 10^{-7}$ $3.08\times 10^{-4}$
Figure 4: Predicted concentration profiles via TumorPred vs reference data obtained from Simcyp Simulator.
4.   Impact and future work
The TumorPred application was initially introduced, validated, and partially utilized in two published papers, [4] and [5]. In [4], the authors developed a novel 9-CNS brain model to measure drug penetration. Subsequently, in [5], they created an R Shiny web application called SpatialCNS-PBPK to make this 9-CNS brain model accessible to a wider audience. To validate the 9-CNS brain model in both studies, the authors compared results from the TumorPred app (which uses a 4-CNS brain model) with those from the Simcyp Simulator v18 [2] (a commercial software which also uses a 4-CNS brain model); this comparison verified the accuracy of TumorPred’s calculations. The authors then packaged the same computational mechanism from the TumorPred app into the SpatialCNS-PBPK app for a 9-CNS model in [4] to evaluate it through concentration-time profiles, sensitivity analysis, and pharmacokinetic parameter (AUC, Cmax, Tmax) calculation. However, TumorPred has an additional feature for estimating drug- and patient-specific parameters, which was beyond the scope of [4] and [5]. TumorPred was later fully utilized in the paper [3] to estimate drug- and patient-specific parameters, drug concentration-time profiles, pharmacokinetic parameters (AUC, Cmax, Tmax), and sensitivity analysis.
TumorPred app provides an invaluable computational modeling tool that improves our ability to prospectively and reliably predict drug penetration and exposure in the human CNS and brain tumors following a successful parameter estimation. The quantitative results obtained through TumorPred, including drug concentration time profiles, Cmax, Tmax, and AUC, can be used to guide the design of efficient clinical trials, selection of right drug candidates, and optimization of dosing regimens to assist rational drug development and treatment for brain cancer. The TumorPred was made easily and freely accessible to the user thus accessible to a wider audience. More complex brain models are already developed for different purposes [6]. Thus, there is a high potential for extending TumorPred for wide range of brain models in future.
Acknowledgements
This research received no funding.